Two types of fibers are suitable for use by a fish net manufacturer. the breaking strength of this fabric is important. it is known that "\sigma"1="\sigma"2=1.0 psi. from random samples of n1=10 and n2=12 we obtain y1=162.5 and y2= 155.0. the company will not adopt fiber 1 unless its breaking strength exceeds that of fiber 2 by at least 10 psi. based on the the sample information, should they use fiber 1? in answering this question, set up and test appropriate hypotheses using "\alpha"=0.01. construct a 99 percent confidence interval on the true mean difference in breaking strength.


1
Expert's answer
2021-06-29T14:37:23-0400

a) The following null and alternative hypotheses need to be tested:

"H_0:\mu_1-\mu_2\geq10"

"H_1:\mu_1-\mu_2<10"

This corresponds to a left-tailed test, for which a z-test for one mean, with known population standard deviation will be used.

Based on the information provided, the significance level is "\alpha=0.01," and the critical value for a left-tailed test is "z_c=--2.3263."

The z-statistic is computed as follows:


"z=\dfrac{\bar{x}_1-\bar{x}_2-\mu}{\sqrt{\dfrac{\sigma^2_1}{n_1}+\dfrac{\sigma_2^2}{n_2}}}"

"=\dfrac{162.5-155-10}{\sqrt{\dfrac{(1)^2}{10}+\dfrac{(1)^2}{12}}}=-5.8387"

Since it is observed that "z=-5.8387<-2.3263=z_c," it is then concluded that the null hypothesis is rejected.

Using the P-value approach: The p-value is "p=P(z<-5.8387)\approx0," and since "p=0<0.01=\alpha," it is concluded that the null hypothesis is rejected.


Therefore, there is not enough evidence to claim that the breaking strength of fabric1 exceeds the breaking strength of fabric 2 by at least 10 psi, at the "\alpha=0.01" significance level.


b) The critical value for "\alpha=0.01" is "z_c=z_{1-\alpha/2}=2.5758."

The corresponding confidence interval is computed as shown below:


"CI=\bigg(\bar{x}_1-\bar{x}_2-z_c\sqrt{\dfrac{\sigma^2_1}{n_1}+\dfrac{\sigma_2^2}{n_2}},"

"\bar{x}_1-\bar{x}_2+z_c\sqrt{\dfrac{\sigma^2_1}{n_1}+\dfrac{\sigma_2^2}{n_2}}\bigg)"

"=\bigg(162.5-155-2.5758\sqrt{\dfrac{(1)^2}{10}+\dfrac{(1)^2}{12}},"

"162.5-155+2.5758\sqrt{\dfrac{(1)^2}{10}+\dfrac{(1)^2}{12}}\bigg)"

"=(6.397, 8.603)"

Therefore, based on the data provided, the 99% confidence interval for the population mean is "6.397<\mu<8.603," which indicates that we are 99% confident that the true population mean "\mu" is contained by the interval "(6.397, 8.603)."